OpenStudy (andpeggy):

The difference between two integers is 13 and their sum is 87. What are the two integers?

OpenStudy (johnweldon1993):

Let the first integer = x and the second = y \[\large x - y = 13\] \[\large x + y = 87\] What would be done next?

OpenStudy (andpeggy):

would you solve by solution/substitution?

OpenStudy (johnweldon1993):

You can either do substitution or elimination *Note that here elimination works much quicker..simply add the 2 equations

OpenStudy (andpeggy):

so using elimination it would turn out being 2x+2y=100 or is that incorrect?

OpenStudy (johnweldon1993):

Not quite When you do: x + x you do indeed get 2x However..what is -y + y ?

OpenStudy (andpeggy):

0

OpenStudy (johnweldon1993):

Exactly! So by adding the 2 equations we have a resulting equation of \[\large 2x = 100\] meaning x = ?

OpenStudy (andpeggy):

x=50

OpenStudy (johnweldon1993):

Good! Anddddd y = ?

OpenStudy (andpeggy):

wasn't it 0?

OpenStudy (johnweldon1993):

it WAS when we combined the equations...however now, go back to the original equations You now know x = 50...so what does 'y' equal in order to satisfy the original equations?

OpenStudy (andpeggy):

is it 1?

OpenStudy (johnweldon1993):

\[\large x - y = 13\] \[\large x + y = 87\] We know x = 50...what does y = ?

OpenStudy (andpeggy):

y=37